OpenStudy (anonymous):

At the deli, sandwiches with 6 ounces of meat cost $3.89. You can order extra meat for $0.25 per ounce. If Greg's sandwich cost him $5.14, how much extra meat,m, did he order? Please put in an equation.

OpenStudy (anonymous):

I didn't have to @ you this time.

OpenStudy (anonymous):

yeah I was watching for it

OpenStudy (anonymous):

Stalker much Jk

OpenStudy (anonymous):

would it not be? 3.89 +0.25x =5.14 then 0.25x = 5.14 -3.89 0.25x = 1.25 x = 5

OpenStudy (anonymous):

exactly what @iTrap said

OpenStudy (anonymous):

Thank you @iTrap